eden made a rectangular mosaic using 256 one centimeter square tiles. she used exactly 80 cm wood to make the frame. what are the dimensions of her design?
Answer by rothauserc(4718)

we have two equations in 2 unknowns
:
let l be length and w be width, then
:
1) 2l + 2w = 80
:
2) l * w = 256
:
solve equation 1 for l
:
l = 40 - w
:
substitute for l in equation 2
:
(40 - w) * w = 256
:
-w^2 +40w -256 = 0
:
(-w +32) * (w -8) = 0
:
w = 32 or w = 8
:
**********************************
:
by convention length > width, so
:
length = 32 cm and width = 8 cm
:
**********************************
: